OpenStudy (nuccioreggie):

For a class pizza party, Esther bought 9 small bottles of juice and ten large bottles of water. Let j represent the cost of juice and w represent the cost of water. Write an expression to determine the total cost for the drinks. Then, calculate the cost if juice costs one dollar and water costs two dollars. ten j plus nine w; twenty-nine dollars ten j plus nine w; twenty-eight dollars nine j plus ten w; twenty-eight dollars nine j plus ten w; twenty-nine dollars

OpenStudy (nuccioreggie):

@Mokeira

OpenStudy (mokeira):

\[9j+10w\]

OpenStudy (nuccioreggie):

o so it will be b in this case?

OpenStudy (mokeira):

and j=1 w=10 so it will be 9(1)+10(10)=9+20=29

OpenStudy (mokeira):

nine j plus ten w; twenty-nine dollars